# The Logical Structure of Physical Laws: A Fixed Point Reconstruction

### v1

@@ Abstract
We formalise the self-referential definition of physical laws (“self-subsumption”) using monotone operators on a lattice of theories, resolving the pathologies of naive set-theoretic formulations. By invoking Tarski’s fixed point theorem, we identify physical theories as least fixed points of admissibility constraints derived from Galois connections. We demonstrate that QED and General Relativity can be represented in such a logical structure with respect to their symmetry and locality principles.
@@ 1 Introduction
A recurring thought in both metaphysics and physics is that lawhood is tied to a kind of stability: Laws are not merely true regularities, but statements (or structures) that remain robust under appropriate changes of description, coordinate choice, or background assumptions. In philosophy, this idea is often expressed by appealing to invariance. What is objective, and what counts as a genuine explanatory pattern, should be invariant under an admissible class of transformations [Nozick2001, Nozick1981]. In physics the same theme appears concretely as symmetry and constraint principles. A mature theory is not just a list of dynamical equations; it is a package that is closed under the consequences of its constitutive requirements (symmetry, locality, unitarity, renormalisability, covariance, etc.).
This paper studies a particularly strong way of packaging that stability idea, which we call self-subsumption, that is, the proposal that the selected laws are exactly those candidates that satisfy an admissibility criterion whose evaluation itself depends on the totality of selected laws. Schematically, one tries to say
@@
Read naively and extensionally, this easily collapses into a Russell-style type confusion. The criterion is applied to an object of the wrong kind, and if one forces well-formedness by collapsing levels, the resulting statement degenerates into a toxic self-membership condition of the form $P\in P$ [Russell1908]. Our first aim is to make this failure precise and to isolate what, in the extensional encoding, turns an intended explanatory constraint into either ill-typed syntax or a trivial self-endorsement.
Our second aim is to give a repair that preserves the self-subsumption idea while preventing the type collapse. The central move is to keep distinct the type of laws (elements of a candidate domain) and the type of law-packages (sets, subspaces, or other structured collections of laws). Once this distinction is enforced, the self-subsumption biconditional becomes a fixed point equation for an induced admissibility operator. In the basic setting where packages form a complete lattice and the admissibility operator is monotone, Tarski’s theorem guarantees fixed points and provides canonical extremal solutions, in particular a least fixed point $\mu\mathcal{F}$ [Tarski1955]. This brings self-reference into a familiar and mathematically controlled form, analogous in spirit (though not identical in purpose) to fixed point treatments of semantic self-reference [Kripke1975].
Our third aim is to explain how this fixed point normal form functions as a useful template in physics. The formalism is not proposed as a machine that derives the laws of nature from first principles. Rather, it isolates a structural pattern already implicit in practice: Once a candidate domain $\Sigma$ is fixed and one specifies admissibility constraints, one obtains a completion map $\mathcal{F}$ whose fixed points are exactly the closed, admissible theory-packages. Importantly, such completion maps need not be inflationary. A ”seed” $S$ may serve only to extract constraints that determine an admissible completion $\mathcal{F}(S)$, without requiring $S\subseteq\mathcal{F}(S)$. This feature matches common physics methodology, where a partial commitment can select a canonical completion under additional principles.
@@ 2 Failure of Extensional Set-Theoretic Formulation
But $P$ was constructed as a subset $P\subseteq\Lambda$, i.e. a law-set (a set of laws, possibly defined under certain criteria) rather than a law. Treating $P$ simultaneously as an element of $\Lambda$ is a type collapse [Russell1908].
@@
In non-well-founded set theories, $P\in P$ can be consistent [Aczel1988]. However, in the extensional encoding, the intended explanation
@@ 3 A Typed Fixed Point Reconstruction
Let $\Lambda$ be a set of Gödel codes of candidate laws (sentences). Let $\Psi:=\mathcal{P}(\Lambda)$ be the complete lattice of law-sets ordered by $\subseteq$. [^4] The powerset $(\mathcal{P}(\Lambda),\subseteq)$ is a complete lattice. For any family $\{S_{i}\}_{i\in I}\subseteq\mathcal{P}(\Lambda)$, arbitrary joins are unions and arbitrary meets are intersections [davey2002]: $\bigvee_{i\in I}S_{i}=\bigcup_{i\in I}S_{i},\qquad\bigwedge_{i\in I}S_{i}=\bigcap_{i\in I}S_{i}.$
@@
Theorem 1 (Tarski’s fixed point theorem [Tarski1955] in specialised form).
@@
where $\mathsf{Obtain}(n)$ is read as ”the law coded by $n$ obtains” (i.e. is among the selected/actualised laws). [^7] We use $\mathsf{Obtain}$ rather than an object-language truth predicate $\mathsf{True}$ to avoid importing a separate theory of truth. A total $\mathsf{True}$ for the same language is delicate in strong settings (Tarski-style undefinability), while Kripke-style approaches treat $\mathsf{True}$ as partial via a fixed point semantics. One could nevertheless replace $\mathsf{Obtain}(n)$ by a suitably interpreted $\mathsf{True}(n)$ (e.g. metalinguistically as satisfaction in the intended model(s), or internally via a Kripke fixed point). This substitution does not affect the fixed point structure developed in the paper, since our main results concern closure of law-sets under the admissibility operator $\mathcal{F}(S)=\{n\in\Lambda:C(S,n)\}$ and the canonical selection $P=\mu\mathcal{F}$, rather than properties specific to truth predicates. [Tarski1955, Kripke1975]
@@
    different fixed points correspond to different admissible global law-sets satisfying the same meta-criterion, [^9] One can conceive a duality à la De Haro and Butterfield [deharo2025duality, deharo2025geometricviewtheories], where any such distinct fixed point characterisation of a theory is isomorphic.
@@
The equation $P=\mathcal{F}(P)$ is still a circular characterisation. The fixed point method does not erase circularity; it replaces it with a non-vicious circularity [^10] Here ”non-vicious circularity” means that self-reference is implemented as a constraint/closure condition rather than as a mere justificatory loop. That is, one specifies a monotone operator $\mathcal{F}$ and then takes the least fixed point $\mu\mathcal{F}$, so the circularity $P=\mathcal{F}(P)$ determines $P$ by minimal closure rather than presupposing it. This is the familiar pattern behind Kripke-style fixed point treatments of self-reference in truth theory, and it aligns with contemporary grounding discussions in which some circles need not be vicious, once a principled notion of ”viciousness” is adopted. [Tarski1955, Kripke1975, Bliss2013] resolved by a selection principle, that is, ”choose the least fixed point $P=\mu\mathcal{F}$”.
@@ 4.1 A General Logical Architecture
$S_{1}\subseteq S_{2}\implies\operatorname{Sym}(S_{2})\subseteq\operatorname{Sym}(S_{1})\implies\operatorname{Inv}(\operatorname{Sym}(S_{1}))\subseteq\operatorname{Inv}(\operatorname{Sym}(S_{2})).$ (1)
@@
$\mathcal{F}(S)\;:=\;\operatorname{Inv}(\operatorname{Sym}(S)).$ (2)
@@ 5 Examples
In the physics examples below, we show that once one fixes a natural candidate domain and encodes standard physical constraints (symmetry, locality, and a truncation criterion) as a monotone completion operator, familiar theories such as QED and GR can be presented as fixed points of that operator. For more information on the subjects, the reader is kindly directed to [Carroll2019, MisnerThorneWheeler1973, PeskinSchroeder2015, Weinberg1995].
@@
    a power-counting bound of dimension $\leq 4$ [PeskinSchroeder2015];
@@
Let $\Gamma$ be a poset of symmetry groups acting on the fields, ordered by inclusion. Since we fix Minkowski spacetime and restrict attention to Lorentz-scalar densities, spacetime symmetries are treated as background constraints and omitted from $\Gamma$ (although it is trivial to incorporate them, as well as the discrete symmetries). [^16] If the spinor field were not massive, it would also admit a transformation of the form $\psi\mapsto e^{i\theta\gamma^{5}}\psi,$ (3) yielding an additional axial $U(1)_{A}$ symmetry at the classical level, so the continuous internal symmetry enlarges to $U(1)_{V}\times U(1)_{A}$ in that limit. Since $\psi$ carries no additional flavour or internal index, the relevant internal symmetry transformations are constant, non-spacetime field redefinitions $\psi\mapsto U\psi$ that preserve the free Dirac density $\bar{\psi}(i\gamma^{\mu}\partial_{\mu}-m)\psi$. Preservation of the mass term requires $\bar{\psi}\psi\mapsto\bar{\psi}\psi$, and preservation of the kinetic term requires $\bar{\psi}\gamma^{\mu}\partial_{\mu}\psi\mapsto\bar{\psi}\gamma^{\mu}\partial_{\mu}\psi$. These two requirements imply that $U$ commutes with the Dirac bilinear structure, hence $U$ can only act as multiplication by a constant phase. Imposing unitarity restricts this to $U=e^{i\theta}$, so the continuous internal symmetry group is $U(1)_{\mathrm{glob}}$. The free theory has a global symmetry, $U(1)_{\mathrm{glob}}$, whereas with the addition of the gauge principle as one of the criteria, we also have a local one, $U(1)_{\mathrm{loc}}$. Therefore,
@@
$\psi(x)\mapsto e^{i\alpha(x)}\psi(x),\;\bar{\psi}(x)\mapsto\bar{\psi}(x)e^{-i\alpha(x)},\;A_{\mu}(x)\mapsto A_{\mu}(x)-\frac{1}{e}\partial_{\mu}\alpha(x),$ (4)
@@
$\mathcal{F}_{\mathrm{QED}}(S)\;:=\;\operatorname{Inv}\!\big(\operatorname{Loc}(\operatorname{Sym}(S))\big)\ \subseteq\ \Sigma_{\leq 4}.$ (5)
@@
    Using (4),
@@
$P_{\mathrm{QED}}=\mathrm{span}\Big\{\big[\bar{\psi}\,i\gamma^{\mu}D_{\mu}\psi\big],\;\big[\bar{\psi}\psi\big],\;\big[-\tfrac{1}{4}F_{\mu\nu}F^{\mu\nu}\big]\Big\}.$ (6)
@@
Let $\Sigma_{\leq 2}$ be the vector space of local scalar densities built from $g_{\mu\nu}$ and finitely many derivatives, restricted to at most two derivatives of the metric, [^18] Restricting the equations of motion to second derivative order is required to avoid the appearance of negative-energy ghost modes, called Ostrogradsky instability. In four dimensions, Lovelock’s theorem [Lovelock1971] guarantees that the only diffeomorphism-invariant actions satisfying this stability criterion are the Einstein-Hilbert and cosmological terms. modulo the standard equivalence relation $\sim$ identifying densities that differ by a total derivative,
@@
$\Gamma:=\{\mathbf{1},\operatorname{Diff}(M)\}.$ (7)
@@
$\mathcal{F}_{\mathrm{GR}}(S)\;:=\;\operatorname{Inv}\!\big(\operatorname{Sym}(S)\big)\ \subseteq\ \Sigma_{\leq 2}.$ (8)
@@
$P_{\mathrm{GR}}:=\mathrm{span}\Big\{\,[\sqrt{-g}],\ [\sqrt{-g}\,R]\,\Big\}\ \subseteq\ \Sigma_{\leq 2}.$ (9)

### v2

@@ Abstract
We formalise the self-referential idea of physical lawhood (“self-subsumption”) as the requirement that a theory contains exactly those candidate laws that are admissible relative to the theory itself. We show that a naive extensional formulation collapses into a Russell-style type confusion (and, if forced, a self-membership pathology). The repair is to distinguish the type of law-candidates from the type of law-packages and to encode admissibility as an operator $\mathcal{F}$ on a lattice of packages; self-subsumption then becomes the fixed-point equation $S=\mathcal{F}(S)$. Under standard assumptions (packages forming a complete lattice and $\mathcal{F}$ monotone), Tarski’s fixed point theorem yields a canonical least fixed point $\mu F$, interpreted as a minimal stable theory under the chosen admissibility constraints. We construct broad classes of such monotone admissibility operators from invariance principles via Galois connections, and we illustrate the schema with toy instantiations inspired by quantum electrodynamics and general relativity that capture symmetry and locality constraints. We do not propose a procedure that derives the laws of nature from first principles; rather, we provide a general fixed-point architecture for reconstructing stable theory-packages from explicit admissibility criteria.
Keywords: laws of nature, logic, symmetry and invariance, philosophy of physics
@@ 1 Introduction
Background and positioning
The contemporary debate on laws of nature spans two broad families of views. On the one hand, Humean and “best system” approaches treat laws as those generalisations (or axioms) that best balance virtues such as simplicity and strength within an overall deductive system [5, 16, 17]. On the other hand, anti-Humean “governing” accounts aim to ground lawhood in features of the world over and above mere regularities, for instance by appealing to relations among universals or other truth-makers for nomological statements [12, 26, 2]. A further line of critique stresses that many celebrated “laws” in physics function via idealisation, ceteris paribus clauses, or model-based approximations rather than strict exceptionless truths [7]. [^2] For a structured overview of these positions and the role of the best-system tradition in particular, see [5]. For an account emphasising the modal and counterfactual robustness of laws and their connection to scientific explanation, see [15].
The present work is intended to be largely orthogonal to these metaphysical commitments. Instead of proposing a new analysis of what laws “are”, we provide a general formal template for law-selection. We represent candidate theories as sets of law-candidates and define an admissibility operator that maps any provisional theory to the set of candidates admissible relative to it. This isolates, in a single mathematical object, the various constraints that are often invoked in practice, including symmetry and invariance considerations [13, 4], while allowing the admissibility notion to depend on the current theory.
Technically, our construction relies on standard fixed-point machinery for monotone operators on complete lattices, where the least fixed point serves as a canonical “minimal stable” theory [24]. This perspective is closely related to the use of closure operators induced by adjoint pairs (Galois connections) and to fixed-point semantics familiar from program analysis [8]. It is also reminiscent, at the level of method rather than subject matter, of fixed-point constructions employed to manage semantic self-reference [14, 25]. Our focus, however, is not the semantics of truth predicates; it is the reconstruction of a stable, constraint-closed theory from an explicit admissibility criterion.
From invariance to self-subsumption
Another recurring thought in both metaphysics and physics is that lawhood is tied to a kind of stability: Laws are not merely true regularities, but statements (or structures) that remain robust under appropriate changes of description, coordinate choice, or background assumptions. In philosophy, this idea is often expressed by appealing to invariance. What is objective, and what counts as a genuine explanatory pattern, should be invariant under an admissible class of transformations [21, 20]. In physics the same theme appears concretely as symmetry and constraint principles. A mature theory is not just a list of dynamical equations; it is a package that is closed under the consequences of its constitutive requirements, such as symmetry, locality, unitarity, renormalisability, covariance, etc.
This paper studies a particularly strong way of packaging that stability idea, which is called self-subsumption, that is, the proposal that the selected laws are exactly those candidates that satisfy an admissibility criterion whose evaluation itself depends on the totality of selected laws. Schematically, one tries to say
@@
Read naively and extensionally, this easily collapses into a Russell-style type confusion. The criterion is applied to an object of the wrong kind, and if one forces well-formedness by collapsing levels, the resulting statement degenerates into a toxic self-membership condition of the form $P\in P$ [23]. Our first aim is to make this failure precise and to isolate what, in the extensional encoding, turns an intended explanatory constraint into either ill-typed syntax or a trivial self-endorsement.
Our second aim is to give a repair that preserves the self-subsumption idea while preventing the type collapse. The central move is to keep distinct the type of laws (elements of a candidate domain) and the type of law-packages (sets, subspaces, or other structured collections of laws). Once this distinction is enforced, the self-subsumption biconditional becomes a fixed point equation for an induced admissibility operator. In the basic setting where packages form a complete lattice and the admissibility operator is monotone, Tarski’s theorem guarantees fixed points and provides canonical extremal solutions, in particular a least fixed point $\mu\mathcal{F}$ [24]. This brings self-reference into a familiar and mathematically controlled form, analogous in spirit (though not identical in purpose) to fixed point treatments of semantic self-reference [14].
Our third aim is to explain how this fixed-point normal form functions as a useful template in physics. The formalism is not proposed as a machine that derives the laws of nature from first principles. Rather, it isolates a structural pattern already implicit in practice: Once a candidate domain $\Sigma$ is fixed and one specifies admissibility constraints, one obtains a completion map $\mathcal{F}$ whose fixed points are exactly the closed, admissible theory-packages. Importantly, such completion maps need not be inflationary. A ”seed” $S$ may serve only to extract constraints that determine an admissible completion $\mathcal{F}(S)$, without requiring $S\subseteq\mathcal{F}(S)$. This feature matches common physics methodology, where a partial commitment can select a canonical completion under additional principles.
@@
We should also note that we do not claim uniqueness or correctness of $\Sigma$ or $\mathcal{F}$; those encode substantive physical choices. Our contribution is the fixed-point architecture and its invariance-based construction of admissibility operators.
@@ 2 Failure of Extensional Set-Theoretic Formulation
But $P$ was constructed as a subset $P\subseteq\Lambda$, i.e. a law-set (a set of laws, possibly defined under certain criteria) rather than a law. Treating $P$ simultaneously as an element of $\Lambda$ is a type collapse [23].
@@
In non-well-founded set theories, $P\in P$ can be consistent [1]. However, in the extensional encoding, the intended explanation
@@ 3 A Typed Fixed Point Reconstruction
Let $\Lambda$ be a set of Gödel codes of candidate laws (sentences). Let $\Psi:=\mathcal{P}(\Lambda)$ be the complete lattice of law-sets ordered by $\subseteq$. [^5] The powerset $(\mathcal{P}(\Lambda),\subseteq)$ is a complete lattice. For any family $\{S_{i}\}_{i\in I}\subseteq\mathcal{P}(\Lambda)$, arbitrary joins are unions and arbitrary meets are intersections [9]: $\bigvee_{i\in I}S_{i}=\bigcup_{i\in I}S_{i},\qquad\bigwedge_{i\in I}S_{i}=\bigcap_{i\in I}S_{i}.$
@@
Theorem 1 (Tarski’s fixed point theorem [24] in specialised form).
@@
where $\mathsf{Obtain}(n)$ is read as ”the law coded by $n$ obtains” (i.e. is among the selected/actualised laws). [^8] We use $\mathsf{Obtain}$ rather than an object-language truth predicate $\mathsf{True}$ to avoid importing a separate theory of truth. A total $\mathsf{True}$ for the same language is delicate in strong settings (Tarski-style undefinability), while Kripke-style approaches treat $\mathsf{True}$ as partial via a fixed point semantics. One could nevertheless replace $\mathsf{Obtain}(n)$ by a suitably interpreted $\mathsf{True}(n)$ (e.g. metalinguistically as satisfaction in the intended model(s), or internally via a Kripke fixed point). This substitution does not affect the fixed point structure developed in the paper, since our main results concern closure of law-sets under the admissibility operator $\mathcal{F}(S)=\{n\in\Lambda:C(S,n)\}$ and the canonical selection $P=\mu\mathcal{F}$, rather than properties specific to truth predicates. [24, 14]
@@
    different fixed points correspond to different admissible global law-sets satisfying the same meta-criterion, [^10] One can conceive a duality à la De Haro and Butterfield [10, 11], where any such distinct fixed point characterisation of a theory is isomorphic.
@@
The equation $P=\mathcal{F}(P)$ is still a circular characterisation. The fixed point method does not erase circularity; it replaces it with a non-vicious circularity [^11] Here ”non-vicious circularity” means that self-reference is implemented as a constraint/closure condition rather than as a mere justificatory loop. That is, one specifies a monotone operator $\mathcal{F}$ and then takes the least fixed point $\mu\mathcal{F}$, so the circularity $P=\mathcal{F}(P)$ determines $P$ by minimal closure rather than presupposing it. This is the familiar pattern behind Kripke-style fixed point treatments of self-reference in truth theory, and it aligns with contemporary grounding discussions in which some circles need not be vicious, once a principled notion of ”viciousness” is adopted. [24, 14, 3] resolved by a selection principle, that is, ”choose the least fixed point $P=\mu\mathcal{F}$”.
@@ 4.1 A General Logical Architecture
$S_{1}\subseteq S_{2}\implies\operatorname{Sym}(S_{2})\subseteq\operatorname{Sym}(S_{1})\implies\operatorname{Inv}(\operatorname{Sym}(S_{1}))\subseteq\operatorname{Inv}(\operatorname{Sym}(S_{2})).$
@@
$\mathcal{F}(S)\;:=\;\operatorname{Inv}(\operatorname{Sym}(S)).$
@@ 5 Examples
In the physics examples below, we show that once one fixes a natural candidate domain and encodes standard physical constraints (symmetry, locality, and a truncation criterion) as a monotone completion operator, familiar theories such as QED and GR can be presented as fixed points of that operator. [^15] Here, for clarity, we should once again note that this approach is constructive; that is, the following is not an argument of derivation. For more information on the subjects, the reader is kindly directed to [6, 19, 22, 27].
@@
    a power-counting bound of dimension $\leq 4$ [22];
@@
Let $\Gamma$ be a poset of symmetry groups acting on the fields, ordered by inclusion. Since we fix Minkowski spacetime and restrict attention to Lorentz-scalar densities, spacetime symmetries are treated as background constraints and omitted from $\Gamma$ (although it is trivial to incorporate them, as well as the discrete symmetries). [^18] If the spinor field were not massive, it would also admit a transformation of the form $\psi\mapsto e^{i\theta\gamma^{5}}\psi,$ yielding an additional axial $U(1)_{A}$ symmetry at the classical level, so the continuous internal symmetry enlarges to $U(1)_{V}\times U(1)_{A}$ in that limit. Since $\psi$ carries no additional flavour or internal index, the relevant internal symmetry transformations are constant, non-spacetime field redefinitions $\psi\mapsto U\psi$ that preserve the free Dirac density $\bar{\psi}(i\gamma^{\mu}\partial_{\mu}-m)\psi$. Preservation of the mass term requires $\bar{\psi}\psi\mapsto\bar{\psi}\psi$, and preservation of the kinetic term requires $\bar{\psi}\gamma^{\mu}\partial_{\mu}\psi\mapsto\bar{\psi}\gamma^{\mu}\partial_{\mu}\psi$. These two requirements imply that $U$ commutes with the Dirac bilinear structure, hence $U$ can only act as multiplication by a constant phase. Imposing unitarity restricts this to $U=e^{i\theta}$, so the continuous internal symmetry group is $U(1)_{\mathrm{glob}}$. The free theory has a global symmetry, $U(1)_{\mathrm{glob}}$, whereas with the addition of the gauge principle as one of the criteria, we also have a local one, $U(1)_{\mathrm{loc}}$. Therefore,
@@
$\psi(x)\mapsto e^{i\alpha(x)}\psi(x),\;\bar{\psi}(x)\mapsto\bar{\psi}(x)e^{-i\alpha(x)},\;A_{\mu}(x)\mapsto A_{\mu}(x)-\frac{1}{e}\partial_{\mu}\alpha(x),$
@@
$\mathcal{F}_{\mathrm{QED}}(S)\;:=\;\operatorname{Inv}\!\big(\operatorname{Loc}(\operatorname{Sym}(S))\big)\ \subseteq\ \Sigma_{\leq 4}.$
@@
    Using the transformation properties of the field and gauge field,
@@
$P_{\mathrm{QED}}=\mathrm{span}\Big\{\big[\bar{\psi}\,i\gamma^{\mu}D_{\mu}\psi\big],\;\big[\bar{\psi}\psi\big],\;\big[-\tfrac{1}{4}F_{\mu\nu}F^{\mu\nu}\big]\Big\}.$
@@
Let $\Sigma_{\leq 2}$ be the vector space of local scalar densities built from $g_{\mu\nu}$ and finitely many derivatives, restricted to at most two derivatives of the metric, [^20] Restricting the equations of motion to second derivative order is required to avoid the appearance of negative-energy ghost modes, called Ostrogradsky instability. In four dimensions, Lovelock’s theorem [18] guarantees that the only diffeomorphism-invariant actions satisfying this stability criterion are the Einstein-Hilbert and cosmological terms. modulo the standard equivalence relation $\sim$ identifying densities that differ by a total derivative,
@@
$\Gamma:=\{\mathbf{1},\operatorname{Diff}(M)\}.$
@@
$\mathcal{F}_{\mathrm{GR}}(S)\;:=\;\operatorname{Inv}\!\big(\operatorname{Sym}(S)\big)\ \subseteq\ \Sigma_{\leq 2}.$
@@
$P_{\mathrm{GR}}:=\mathrm{span}\Big\{\,[\sqrt{-g}],\ [\sqrt{-g}\,R]\,\Big\}\ \subseteq\ \Sigma_{\leq 2}.$